### v2

# Pilot-Quantum: A Quantum-HPC Middleware for Resource, Workload and Task Management

Abstract
As quantum hardware continues to scale, managing the heterogeneity of resources and applications – spanning diverse quantum and classical hardware and software frameworks – becomes increasingly critical. Pilot-Quantum addresses these challenges as a middleware designed to provide unified application-level management of resources and workloads across hybrid quantum-classical environments. It is built on a rigorous analysis of existing quantum middleware systems and application execution patterns. It implements the Pilot Abstraction conceptual model, originally developed for HPC, to manage resources, workloads, and tasks. It is designed for quantum applications that rely on task parallelism, including: (i) Hybrid algorithms, such as variational approaches, and (ii) Circuit cutting systems, used to partition and execute large quantum circuits. Pilot-Quantum facilitates seamless integration of quantum processing units (QPUs), classical CPUs, and GPUs, while supporting high-level programming frameworks like Qiskit and Pennylane. This enables users to design and execute hybrid workflows across diverse computing resources efficiently. The capabilities of Pilot-Quantum are demonstrated through mini-applications – simplified yet representative kernels focusing on critical performance bottlenecks. We present several mini-apps, including circuit execution across hardware and simulator platforms (e. g., IBM’s Eagle QPU), distributed state vector simulation, circuit cutting, and quantum machine learning workflows, demonstrating significant scale (e. g., a 41-qubit simulation on 256 GPUs) and speedups (e. g., 15x for QML, 3.5x for circuit cutting).
Index Terms:
Quantum Computing, HPC, Middleware
I Introduction
Motivation: Quantum computing promises advancements in computational challenges across science and industry, with applications in chemistry, simulations, optimization, and machine learning [1]. Quantum algorithms leverage quantum information principles to outperform classical methods with fewer computational steps [2, 3]. While Noisy Intermediate-Scale Quantum (NISQ) hardware [4] has progressed, further improvements in qubit fidelity [5], error correction [6], and software are needed for a practical advantage [7].
Quantum-HPC Integration: Quantum computing complements HPC by addressing computational challenges currently intractable. HPC supports quantum research by enabling scalable classical simulations essential for many hybrid quantum-classical algorithms. Tighter integration of classical and quantum resources will become increasingly important [8, 9, 10, 11, 12]. This integration presents challenges across several levels: (i) Hardware Integration: Efficiently combining heterogeneous quantum systems (e.g., superconducting and ion-trap qubits) with classical computing resources (e. g., GPUs and CPUs) to create cohesive hybrid systems. (ii) Software Ecosystems: Developing robust software frameworks, tools, and middleware systems that seamlessly integrate with existing HPC tools (e. g., CUDA and MPI runtimes). Achieving this integration is crucial for advancing quantum computing and unlocking its full potential in solving complex computational problems.
Challenges: In particular, software tools [13, 14, 15] and middleware [16, 17, 18] will become critical for hardware development. The fragmented quantum software stack, still in its infancy, requires integration across multiple layers – from qubit control and circuit execution to application workflows. NISQ and FTQC algorithms rely heavily on quantum-classical integration [19, 20]. With maturing algorithms, end-to-end application workflows integrating data acquisition, pre-processing, and post-processing will gain importance [21]. Resource, workload, and task management are critical as these workflows use heterogeneous resources. Coordinating quantum and classical resources, managing tasks across diverse infrastructures, and integrating multiple programming paradigms complicate quantum application scaling.
Quantum-HPC Middleware: We propose Pilot-Quantum, a middleware for resource, workload, and task management in quantum-HPC systems. It is based on the Pilot-Abstraction [22, 23] and a conceptual quantum-HPC middleware architecture [20], designed from an analysis of quantum-HPC execution patterns (motifs [19]). The pilot abstraction uses placeholder jobs to reserve and manage resources across heterogeneous infrastructures (e. g., HPC and clouds). It supports loosely coupled parallelism in quantum applications, e. g., variational algorithms and circuit cutting.
Pilot-Quantum enables resource management and scheduling at the application level. It integrates seamlessly with quantum software ecosystems, supporting both high-level frameworks like Qiskit and Pennylane and low-level frameworks like cuQuantum for hybrid workloads [19]. With application-level scheduling, applications can optimize resource allocation, essential for costly QPUs, enabling efficient resource use for diverse classical and quantum task-based workloads.
This paper is structured as follows: We explore quantum software and middleware state-of-the-art in section II. The primary contribution is Pilot-Quantum, an open-source middleware addressing resource, workload, and task management challenges is introduced in section III. Section IV demonstrates its capabilities by extending quantum mini apps [19] to new motifs, including distributed state vector simulation, circuit cutting, and quantum machine learning (QML). We conclude with a qualitative comparison of quantum middleware systems with Pilot-Quantum and future work in section V.
II Background and Related Work
This section analyzes quantum-classical software and middleware using a conceptual architecture model [20] and quantum-HPC integration modes and execution motifs [19].
II-A Conceptual Middleware Architecture
The conceptual quantum-HPC middleware architecture [20] consists of four distinct layers, denoted as $L1$, $L2$, $L3$, $L4$: $L4$ is the workflow layer, which encapsulates application semantics and logical dependencies between the workflow stages and offers a high-level abstraction for developing the quantum-HPC workflows. The workload layer $L3$ manages the workload, a set of interdependent, quantum, and classical tasks emitted by $L4$. It optimizes the workload execution by partitioning tasks and assigning them to the available resources utilizing parallelism (e.g., ensemble and task parallelism).
The $L2$ task layer performs the operational management of task execution on a single resource. Any runtime operations, i.e., task scheduling, assignment, execution, and monitoring, are managed at $L2$. Finally, the resource layer $L1$ describes how a system abstracts and manages heterogeneous classical (e. g., HPC and cloud systems, including accelerators) and quantum resources (e. g., gate-based and analog QPUs). $L1$ layer provides the capability to exploit system-level tools, e. g., GPU acceleration and QPU-target optimization are essential for efficient quantum-classical integration.
II-B Quantum-HPC Integration Modes and Motifs
Managing quantum-HPC workloads and determining quantum-classical task dependencies requires identifying integration modes and motifs (i. e., recurring execution patterns) between quantum-HPC components [19].
Integration modes
are categorized as Quantum-x-HPC or HPC-x-Quantum based on the strength of coupling between quantum and classical tasks [8, 20]. Three modes are identified: HPC-for-Quantum, with strong coupling optimizing quantum-classical interactions (e. g., error mitigation [24], error correction [25, 26], dynamic circuits [27]); Quantum-in-HPC emphasizes loosely-coupled parallelism between classical and quantum resource found in variational quantum algorithms (VQA) [28] to accelerate HPC application; and Quantum-about-HPC, loosely coupling quantum elements into HPC applications for pre- and post-processing [29, 30].
Motifs
[19] are recurring high-level execution patterns describing quantum-classical task interactions, e. g., whether operations occur within the quantum system’s coherence time (for tightly coupled tasks) or beyond it (for loosely coupled tasks). For classical simulation of quantum systems, both coupling components exist. Distributed state vector simulation involves tightly coupled classical tasks using MPI and GPU/CUDA parallelism, while variational algorithms and circuit cutting typically require loose coupling between tasks.
II-C Quantum Software Layers
This section details the quantum software stack, mapping it to the layers of the conceptual model (see Figure 1 and [13]).
Figure 1: Quantum Software Stack: An overview of key components and layers, from quantum programming environments and hybrid runtimes to hardware resource management.
Quantum Programming
frameworks typically provide quantum circuit model-based abstraction for describing quantum programs. Examples are Qiskit [14, 31], PennyLane [15], Cirq [32], low-level frameworks like CUDA-Q and AWS Braket [33], and high-level programming libraries, such as, Qrisp [34], Silq [35], and Q# [36]. At the $L4+$ levels, domain-specific programming libraries include Qiskit Optimization [37], Qiskit Nature [38], and OpenFermion [39].
Hybrid Runtimes
like Qiskit Runtime [40] and Braket Hybrid Jobs [41, 42] provide session and decorator abstractions, respectively, to co-allocate heterogeneous resources for hybrid quantum-classical applications. However, new execution patterns, such as the combination of quantum and machine learning capabilities, require new abstractions, including auto differentiation of quantum circuits (e. g., using PennyLane [15]) or training classical models using quantum task outputs [43]. Some frameworks like XACC [44], QCOR [45], and CUDA-Q [18] use a kernel-based programming model with compile toolchains to manage quantum-classical function interactions. For classical simulation of quantum systems, acceleration frameworks, e. g., cuQuantum [46] are critical for performance and scale.
Hardware Integration, Transpiler/Compiler, and Optimization:
At the $L1$ layer, many quantum frameworks provide device abstractions and plugins for accessing quantum devices, such as Qiskit backends [47], Qiskit Runtime [40], and PennyLane plugins [48]. Intermediate representations like OpenQASM [49] and QIR [50, 51] aim to bridge the gap between hardware platforms and high-level programming constructs. Transpiler and compiler are used to transform a circuit to a hardware-optimized form and are provided by full-stack frameworks like Qiskit and specialized frameworks like Tket [52], BQSKit [53], and Staq [54] transpilers.
Multi-level Resource Management:
Resource management and scheduling in quantum-HPC environments occur across multiple levels and time scales [9, 55], involving integration of coarse-grained job-level allocation (e. g., GPUs) with QPU allocations and fine-grained, application-level scheduling [23, 56]. Existing HPC resource management frameworks often assume exclusive resource access, leading to resource fragmentation and underutilization [57]. As algorithms scale across heterogeneous resources (e. g., CPUs, GPUs, QPUs), such resource management will result in suboptimal performance and utilization. Pilot-Quantum specifically addresses the need for an application-level resource management balancing system and application objectives.
Middleware and System-level Runtimes
at the task and workload layer ($L3$ and $L4$) are needed to orchestrate quantum workloads across integrated quantum-HPC systems. QML workloads, for instance, involve heterogeneous classical-quantum tasks for gradient computation, backpropagation, optimization, and circuit execution, hence requiring a systematic approach to workload and task management.
II-D Middleware for Resource, Workload and Task Management
Many quantum middleware systems emerged recently including, QFaaS [58], Tierkreis [59], Qibo [60], Quantum Framework (QFw) [61], Quantum Brilliance [62], Rasqal [63] and Zapata’s Orquestra [64]. We particularly investigate Qiskit Serverless [16], Covalent [17], and CUDA-Q [18] systems system, which have different design objectives and principles.
II-D1 Qiskit Serverless
[65, 66, 16, 67] serves as middleware for Qiskit, managing resources and workloads in hybrid quantum-classical infrastructures while supporting long-running applications and scaling across distributed resources (CPU, GPU, QPU). It extends cloud-based quantum hardware access to classical computing environments at the $L2$ and $L3$ layers, utilizing Apache Ray [68] to execute hybrid workloads.
Qiskit Serverless maps to the $L2$/$L3$ layer of the conceptual architecture, facilitating loosely and tightly coupled interactions. For workload management at the $L3$ layers, it offers a distributed task decorator for manual task-resource mapping. It interfaces with IBM Quantum Cloud services and hardware providers like IonQ and Braket [69, 70] through Qiskit Runtime and backend APIs. QPU priority is available for computations within the same Qiskit Runtime $Session$, allowing for the co-location of quantum and classical tasks.
II-D2 Covalent
[17, 71] orchestrates distributed computing workflows across Cloud-HPC resources. Its SDK provides $L4$ level abstraction using electron and lattice decorators for defining tasks and workflows (i. e., DAG of electrons), where quantum tasks are defined using qelectron decorator. The Dispatch service at the $L3$ layer accepts workloads (i. e., ready-to-run tasks), maps them to suitable resources, and dispatches them to appropriate backends.
At the $L2$ and $L1$ layers, Covalent uses executors to abstract resources, e. g., so-called QExecutors for quantum resources. Covalent allows grouping executors in structures like QCluster (i. e., set of quantum resources), enabling parallelism within resources managed by one executor (e.g., GPU acceleration) and across multiple executors. Covalent uses Dask [72] as the default task manager and necessitates manual resource configuration, including AWS Dask cluster setup via Terraform, which lacks HPC infrastructure support. Limited HPC support, e. g., via a pilot plugin [73] for Radical-Pilot/PSI/J, exists. For application-level task-to-executor mapping, Covalent’s dispatch service considers electron decorator-specified requirements and executor availability for dynamic task assignment at runtime. However, limited quantum cloud provider support by Covalent prevents co-scheduling of classical and quantum electrons, hindering QPU prioritization.
II-D3 CUDA-Q
[18] supports tight integrations within heterogeneous computing environments. It employs an LLVM compiler toolchain [74] to convert frontend code in C++ or Python to different heterogeneous backends, such as CPUs, CUDA-enabled GPUs, and QPUs. It utilizes distinct intermediate representations for each processor type (e. g., QIR for QPUs and NVVM IR for GPUs), allowing for targeted optimizations. The CUDA-Q runtime, utilizes MPI for parallel processing and CUDA as a low-level runtime. While particularly suitable for HPC-for-Quantum tight integration, CUDA-Q also supports Quantum-for-HPC modes, e. g., for VQAs. At $L4$, it provides a circuit-based programming model and a kernel API for offloading circuits to hardware or simulated QPUs. Application-level task management uses an MPI and CUDA-Runtime on the $L2$/$L1$ level.
III Pilot-Quantum Design and Architecture
Pilot-Quantum enhances resource, workload, and task management across heterogeneous classical and quantum infrastructures for quantum applications and workflows. It is an implementation of the Pilot Abstraction [22, 23], a conceptual abstraction for application-level scheduling [22, 75]. It uses a placeholder job concept for application-level control of allocated resources, crucial for scaling heterogeneous workloads.
Abstraction for Resource Management: Pilot-Quantum’s abstractions enable applications to effectively decompose workflows, simplify scaling, and manage diverse task-level parallelism. It can efficiently parallelize and execute operations like circuit ensembles or data-parallel QML training asynchronously and concurrently.
Easy Deployment: Pilot-Quantum simplifies workflow deployment across distributed, heterogeneous infrastructures, including non-local quantum cloud environments, HPC MPI/CUDA environments for distributed state vector simulation, and loosely coupled workflow systems. It can also integrate with resource-specific optimization tools, such as QPU-specific transpilers.
Application Performance: Pilot-Quantum uses hierarchical, multi-level resource management to align system- and application-level requirements, enabling application-level optimization of workload and resource allocation and use. This is crucial for distributed environments with application demands and varying resource availabilities. For instance, in congested QPU hardware, its multi-level scheduling balances system utilization with application workload execution requirements.
III-A Requirements and Design Objective
Infrastructure Heterogeneity: The system must address both software and hardware heterogeneity, encompassing a range of QPUs such as superconducting qubits, ion traps, neutral atoms, and simulated QPUs utilizing MPI or CUDA. This requires compatibility with diverse frameworks and tools, including transpilers, compilers, and runtimes. Examples include various simulator types (e. g., accelerated state vector and tensor network simulations leveraging MPI or CUDA) and machine learning runtimes like PyTorch and JAX.
Applications: Quantum-HPC applications exhibit diverse workloads [19, 76], as they evolve from monolithic applications to complex multi-stage workflows [77]. It can involve quantum data encoding (e. g., amplitude encoding, Jordan-Wigner transformation for molecules), quantum mechanical system simulation (e. g., ground state energy), and machine learning-based error mitigation strategies. Such workflows require diverse infrastructures and resources across execution stages, whereas classical computation may require accelerators like GPUs for simulation and machine learning.
Resource management in quantum-classical computing is challenging due to diverse resource types and application demands. Dynamic resource allocation during workflow execution is crucial to address application and infrastructure dynamism when handling varied tasks and dependencies across heterogeneous platforms. This requires employing multi-level scheduling coordination with efficient resource allocation, workload partitioning, and task management strategies to further optimize application performance and system utilization.
Abstractions: Integrating quantum computing into HPC workflows requires user-friendly, flexible abstractions that hide quantum complexity while supporting various parallelism forms (e.g., task parallelism for circuits, distributed state vector simulations using MPI and cuQuantum, and differentiable computing for QML). These should facilitate efficient pipeline creation for end-to-end quantum applications with pre- and post-processing integrations. The abstractions must offer seamless interoperability with quantum and HPC infrastructures, further ensuring compatibility with frameworks like CUDA for GPU-accelerated tasks and MPI for distributed computing.
Interoperability: Rapidly evolving quantum programming frameworks across all technology stack layers necessitate low-overhead integration with existing ones at both application and system levels. For instance, it’s crucial to integrate with vendor-specific low-level runtimes that enable tight coupling between quantum and classical resources, such as CUDA-Q’s MPI/CUDA runtime or Qiskit Runtime.
III-B System Architecture
Figure 2: Pilot Quantum Architecture: The system core is a Pilot-Manager that orchestrates and manages resources through pilots across both classical and quantum infrastructures, such as QPUs, GPUs, and CPUs. Pilots are responsible for reserving resources and managing task execution.
Figure 2 shows Pilot-Quantum’s architecture. The Pilot-Job system [22, 23] acts as an intermediary between applications, higher-level systems (e.g., workflow management), and local resource management systems (e. g., Slurm on HPC). It consists of a Pilot-Manager and multiple Pilots. The Pilot-Manager orchestrates resources by provisioning, managing, and releasing them for workload execution. It optimizes task assignment across pilots for efficient time-to-completion and system utilization. The Pilot-Manager provides extensible resource plugins to integrate various resources.
The Pilot Abstraction’s multi-level scheduling suits quantum-HPC environments for two reasons: (i) heterogeneity and complexity of quantum and classical applications complicate static workload and resource assignment, and (ii) early quantum era co-design of applications and hardware requires integrated application-level resource management. While small applications can use hard-wired integration, larger hybrid systems need more adaptive resource management. Hence, hierarchical resource management capabilities are crucial for developing and operating quantum-enabled workflow systems efficiently. Next, we present each layer of the system.
III-B1 L4 Workflow Layer and Abstractions:
Pilot-Quantum provides two abstractions: (i) resource management, enabling applications to allocate appropriate classical and quantum resources, and (ii) workload and task management, facilitating efficient workload execution on these resources. For (i), applications must define required resources in a pilot_description and submit it to Pilot-Quantum, which then acquires the resources.
For (ii), applications must decompose their problems into a set of independently executable tasks, collectively forming the workload. Tasks can encapsulate units of quantum computations (e. g., circuit executions) and classical computations (e. g., pre- and post-processing). While early quantum applications require high control over problem decomposition and workload, Pilot-Quantum is expected to interface with higher-level workflow systems and application libraries.
Tasks are defined using a task_description and submitted via the submit_task API. Pilot-Quantum also integrates with native abstractions of different frameworks, e. g., Dask and Ray. Tasks can be assigned: (i) directly to a specific pilot or (ii) to a group of pilots. For (ii), the pilot manager assigns tasks to pilots, optimizing throughput by load balancing across all pilots. Please, refer to [78] for API details.
III-B2 L3/L2 Workload and Task Layers:
After receiving the pilot_description, the Pilot-Manager allocates requested resources via the local resource management system, (steps 1-3 in Figure 2). In HPC environments, a placeholder job, which starts the Pilot-Agent is queued via Slurm. Once active, the Pilot Agent manages the allocated resources and executes tasks either directly or through external execution engines. External execution engines are supported via plugins; currently, Pilot-Quantum supports Dask [79] and Ray [68]. Further, it integrates with low-level runtimes (e. g., CUDA and MPI). Quantum hardware access is handled via classical tasks that utilize, e. g., Qiskit Providers or PennyLane plugins.
The choice of an external execution engine depends on workload characteristics and resource infrastructure. Dask is well-suited for data-intensive workloads, such as data-parallel streaming and machine learning tasks [79]. Pilot-Quantum extends its capabilities with a plugin for Ray [68], a versatile distributed computing framework that facilitates seamless integration of GPUs (essential for quantum simulations) and custom resources like QPUs. For tightly coupled HPC tasks, such as distributed state vector simulations, the MPI and CUDA launch mechanisms are crucial. The agent leverages Slurm’s srun command to efficiently distribute tasks across nodes, cores, and GPUs, ensuring optimal parallel execution considering machine architecture and network topology.
III-B3 L1 Resource Layer:
As described, Pilot-Quantum aims to abstract heterogeneous resources through its plugin mechanism for various runtimes. In addition, it can be integrated with various resource-specific tools, e. g., QPU-specific transpilers.
IV Evaluation and Results
This section characterizes Pilot-Quantum’s capabilities using the Quantum Mini App framework [19, 76]. Mini-apps are simplified versions of quantum-HPC applications, representing key characteristics and performance bottlenecks through recurring execution patterns (motifs). We extend the circuit simulation mini-app [76] and develop a QML mini-app.
[…]
IV-A Pilot-Quantum Characterization
[…]
$256$ $0.6$ $418.1$
$512$ $1.3$ $406.5$
$1024$ $2.6$ $397.6$
$2048$ $5.8$ $358.6$
$4096$ $11.4$ $358.4$
$8192$ $27.4$ $298.2$
TABLE I: Average Throughput and Duration by Task Size
[…]
IV-B Circuit Simulation Mini-App
We extend the circuit simulation mini-app [19] with quantum hardware support (section IV-B1), an implementation of distributed state vector simulation (section IV-B2), and a circuit cutting motif (section IV-B3).
IV-B1 Circuit Execution:
[…]
Figure 3: Circuit Execution on IonQ, IBM Eagle and Qiskit Aer (CPU and GPU): Comparing the execution times for 1,024 random quantum circuits (2 to 28 qubits) on IonQ Quantum Cloud and Qiskit’s Aer simulators (CPU and GPU) using Pilot-Quantum (Ray) on one Perlmutter node. For IBM Eagle, we only executed eight random circuits for each qubit configuration and scaled their mean execution time to be comparable with the 1,024-circuit experiments. IonQ required 4,260 seconds for 28 qubits, while Aer (CPU) took 469.7 seconds, and Aer (GPU) completed it in just 49 seconds. The scaled execution time of the IBM Eagle QPU is 199 seconds.
We evaluate various hardware and simulated QPU backends, specifically: (i) in-process simulators running on application-allocated resources, like Qiskit Aer [81], (ii) cloud simulators, and (iii) a QPU via the IBM cloud. For (i), we investigate Qiskit’s Aer CPU and GPU simulators. For (ii), we investigate IonQ’s cloud simulator [82]. Further, we investigate IBM’s Eagle QPU hardware accessed via the cloud. We used Pilot-Quantum (Ray) on one Perlmutter node to manage the execution of $1024$ random quantum circuits with $2$-$28$ qubits across these backends.
[…]
For IBM Eagle, we executed only eight random circuits for each qubit configuration and scaled their mean execution time to be comparable with the $1,024$-circuit experiments. In contrast to the simulation experiments, the execution time varies slightly but does not increase exponentially.
IV-B2 Distributed State Vector Simulation:
enables efficient simulation of large quantum circuits with higher qubit counts and depth by using multiple nodes and GPUs [83, 84]. The distributed state vector motif [19] uses PennyLane’s lightning.gpu [85] v0.35.0. This device is a quantum circuit simulator implemented on NVIDIA’s cuQuantum [46], using a C++ backend and CUDA/MPI for distributed computations and data movements.
[…]
Figure 4: PennyLane lightning.gpu Distributed State Vector Simulations: Computing the expectation value of a 2-layer strongly entangling layered (SEL) circuit with and without gradient calculation using PennyLane’s lightning.gpu device on 256 A100 80 GB GPUs on Perlmutter.
Figure 4 shows distributed state vector simulation of a 2-layer strongly entangling layered (SEL) circuit [86]. We compare circuit execution with and without gradient computation, using adjoint differentiation [87] for gradients, which offers lower memory and computational overhead than alternatives like finite difference requiring multiple circuit evaluations. We investigated $31$-$37$ qubits on 256 GPUs, the upper qubit range limited by gradient computation overheads, without which can scale to $41$ qubits. Results indicate GPU resources must be balanced between simulation and other intensive tasks like ML for optimal performance.
IV-B3 Circuit Cutting
[88] enables running larger circuits on quantum hardware at the cost of increased sampling and classical processing overhead. It decomposes larger circuits into smaller ones, referred to as subcircuits, by cutting wires and gates, allowing execution within hardware and simulator constraints. However, the number of required subcircuit executions increases exponentially with circuit size and number of cuts [89]. The sampling results from the subcircuits are combined to reconstruct the original circuit outcome.
This circuit-cutting motif uses Qiskit’s circuit cutting addon [90]. It uses Qiskit’s EfficientSU2 ansatz, commonly used in many variational algorithms, comprising single qubit rotations and linear CNOT gates between adjacent qubits. Qiskit’s expectation value reconstruction combines results from smaller sub-experiments. It provides options to run full circuit and/or circuit cutting simulations, enabling comparison of the error induced by circuit cutting. It gathers comprehensive performance metrics throughout the process.
12481610101010100100100100200200200200Number of NodesTime (seconds)(iii) Full Circuit (Distributed State Vector with MPI/CUDA)(ii) Circuit Cutting (4 cuts, 2100 subexperiments with 12 qubits)(i) Circuit Cutting (2 cuts, 72 subexperiments with 17 qubits)
Figure 5: Qiskit Circuit Cutting technique with Pilot-Quantum: Illustrates performance of circuit cutting for a 34 qubit circuit using 2 cuts/72 subexperiments, 4 cuts/2100 subexperiments, and full circuit simulation using Pilot-Quantum/Ray on Perlmutter 4x A100 80 GB GPUs nodes and Qiskit’s AerSimulator (with CUDA/MPI support). For circuit cutting, we use one GPU per subexperiment. For full circuit simulation, we use 4 GPUs/node and distributed state vector simulation.
In the following experiments, we use a $34$-qubit quantum circuit of the described ansatz. For circuit cutting, we utilize two configurations: (i) subcircuit size $17$ and (ii) subcircuit size $12$. We used the find_cuts method to determine the best cut and generate necessary subcircuit executions: $72$ for scenario (i), and approximately 2,100 subcircuit experiments for (ii). We compare these circuit-cutting scenarios to a full circuit simulation using Qiskit’s Aer distributed state vector simulator (scenario (iii)) with multi-node GPU and MPI support [81, 91]. All experiments are repeated at least three times.
Figure 5 illustrates performance scaling from $1$ to $16$ nodes ($4$ to $64$ GPUs). In scenario (i), with fewer subexperiments, optimal performance is achieved on a single node. Conversely, scenario (ii) shows the best performance with $2$ nodes, achieving a $\sim 1.5$ speedup, demonstrating that additional resources executed the 2,100 subexperiments faster. When comparing to the sequential runtimes on 1 core for (i) 58 sec and (ii) 304 sec, we were able to achieve a (i) 2.1×2.1$\times$2.1 × and (ii) 3.5×3.5$\times$3.5 × speedup. Scenario (iii), the distributed state vector simulation, outperforms both circuit-cutting scenarios when using more than $4$ nodes. The best performance was achieved with 16 nodes, showing a $\times5$ speedup.
In summary, different circuit-cutting strategies lead to varying optimal performance points, highlighting the importance of application-level scheduling, as enabled by Pilot-Quantum, for efficient resource and task management. Hence, an adaptive execution strategy combining multiple simulation methods (e. g., circuit cutting and distributed state vector simulation) and leveraging heterogeneous resources (e. g., CPUs, GPUs, and QPUs) will be crucial for effective application scaling.
IV-C QML Workflow
Quantum Machine Learning (QML) is a key workload [92] enabling error mitigation and new applications [20]. QML workflows typically include data preparation, training, and evaluation, leveraging parallelism for tasks like hyperparameter optimization and architecture searches. Data parallelism processes training subsets simultaneously, while tensor parallelism ensures efficient gradient computation and backpropagation. Future workflows will likely integrate QML models for inference with other tasks. This section presents a QML mini-app with a multi-stage pipeline for data encoding, compression, and classifier training using Pilot-Quantum.
12481632005,00050005{,}0005 , $00010{,}$ $00015{,}$ $00020{,}$ $00025{,}$  [^0] Number of NodesTime (seconds)SweepBFGS0020202020404040406060606080808080100100100100Efficiency (%)Efficiency
Figure 6: Execution time and efficiency for compressing the 60,000 samples of the CIFAR-10 data set for cluster configurations with varying numbers of nodes. The time is divided between the two phases, with sweeping represented by the red portion and BFGS by the blue. The results demonstrate that as the number of nodes increases, the overall execution time decreases while the efficiency reduces slightly, particularly for larger node counts.
IV-C1 Data Encoding and Compression:
Efficiently representing classical data (e.g., images) for QML is critical due to quantum state preparation bottlenecks [93, 94]. Tensor networks, like matrix product states (MPS), can efficiently approximate images [95]. Our QML pipeline mini-app uses Pilot-Quantum (Ray) for a two-stage data compression process. We evaluated the pipeline using CIFAR-10 [96] comprising $10$ classes and $60,000$ images, encoding each $32\times 32$ pixel image with $3$ RGB color channels into a 13-qubit quantum state ($10$ qubits encoding spatial information, and $3$ encoding color channels [97]). The image’s quantum state is approximated by a depth $4$ circuit with 205 trainable parameters, using only single and two-qubit gates on neighboring circuits. The first stage trains the circuit using a sweeping algorithm [98], while the second stage employs parameterized circuit training with a state vector simulator and BFGS optimizer [99].
[…]
Figure 7: Batch processing time by batch size for a variational quantum classifier implemented in PennyLane with JAX. The figure compares four processing methods: sequential (no optimizations), batch optimization using vmap, JIT compilation, and a combination of both. JIT compilation improves runtime by about two orders of magnitude. Batch optimization using vmap decouples the execution time from batch size. The combination of both vmap and JIT provides the best performance, significantly reducing processing time across all batch sizes.
[…]
IV-C2 Classification:
The compressed dataset was classified using a PennyLane-implemented variational quantum classifier on one NVIDIA $A100$ GPU. The circuit uses $13$ qubits, with one arbitrary rotation layer per qubit and a ring of CNOTs [100]. Next, 10 Pauli-Z expectation values generate the prediction. The ADAM optimizer trains the model by minimizing cross-entropy loss between predicted output and target labels. We used PennyLane’s default.qubit with JAX [101] interface, utilizing JIT and vmap for performance optimization.
Figure 7 compares batch processing times for the variational quantum classifier using four methods: sequential (unoptimized), vmap batch optimization, JIT compilation, and combined vmap-JIT. JIT compilation improves runtime by about two orders of magnitude, while vmap decouples execution time from batch size. The vmap-JIT combination performs best, significantly reducing processing time across all batch sizes. This QML mini-app demonstrates how parallelism through Pilot-Quantum (Ray) and optimization tools like JAX can greatly enhance QML workflow efficiency and scalability.
V Discussion and Conclusion
TABLE II: Comparison of quantum-HPC middleware systems
Middleware Qiskit Serverless Covalent CUDA-Q Pilot-Quantum
Characteristics L4 Workflow Layer
Abstraction Task Lattice & electrons Kernel Task
abstraction abstraction (DAG) abstraction abstraction
Quantum Qiskit  Qiskit, CUDA-Q  Qiskit,
libraries  PennyLane, etc.  PennyLane, etc.
L3 Workload Layer
Resource Serverless (cloud) Quantum, HPC, Local resource Pilot-Job
management or container Cloud Executors management
(self-managed) or serverless (e. g., SLURM)
Task Application Application Automatic Application
partitioning level task level task partitioning (only level task
partitioning partitioning quantum tasks) partitioning
L2 Task Layer
Task distributed_- electron None Pilot-Job, Ray
abstraction task or Dask task
Quantum task same as classic QElectrons  cudaq.kernel same as classic
Task Ray Dask CUDA-Q Ray, Dask
management
L1 Resource Layer
Computing Cloud, Docker, HPC, Cloud, HPC (Nvidia HPC, Cloud
Infrastructure Kubernetes (only Kubernetes GPUs)/ Cloud
cores, no GPUs) (Docker-based)
Quantum IBM Quantum, any CUDA-Q any
Infrastructure  Qiskit API (e. g., integrations
IonQ, Rigetti) (e. g., IQM, IonQ)
Runtime Qiskit Runtime any MPI, CUDA any
Simulator Qiskit Aer any cuQuantum any
backends (e. g., (cuStateVec,
cuQuantum) cuTensor)
V-A Discussion
This section qualitatively compares Pilot-Quantum with Qiskit Serverless, Covalent, and CUDA-Q (see section  II-D) by mapping each across the four architectural layers (see Figure 8 and Table II).
L4: Qiskit Serverless and Pilot-Quantum use task-based models, while Covalent provides workflow-level abstraction at $L4$ layer. CUDA-Q’s kernel-based approach supports tighter integration modes (e. g., error correction) and large-scale computational demands (e. g., distributed state vector simulations), and can be used in conjunction with Pilot-Quantum.
While each framework supports workflow dynamism, i. e., enabling new task creation based on prior results, including quantum tasks, only Covalent and Pilot-Quantum allow adding resources at runtime. At the application-level, Qiskit Serverless is designed to integrate with Qiskit’s libraries, while Covalent and Pilot-Quantum are framework-agnostic. In contrast, CUDA-Q is optimized for tight coupling of quantum and classical tasks.
L3: Both Covalent and Pilot-Quantum offer high-level control over resource allocation and task management. Qiskit Serverless allows application-level task management but not resource allocation, while CUDA-Q, uses a fixed HPC resource model with MPI and CUDA for task execution, lacking resource management capabilities and requiring manual integration with resource management systems like Slurm.
L2: At the task execution layer, most frameworks use task-based systems like Dask and Ray for parallel execution across diverse resources, decoupling application from execution details (e. g., scheduling). They allow some application-level control of task and resource mapping for optimally utilizing accelerator and data parallelism. Resource mapping depends heavily on application details, such as the amount of data per task, making application-level scheduling crucial for optimal performance. CUDA-Q, with its runtime, directly manages kernel offloading to QPUs or accelerators. While supporting some dynamism, CUDA-Q’s compiled nature makes it less flexible than other task execution systems.
Figure 8: Comparison of quantum-HPC Middleware: Covalent, Qiskit-Serverless, and Pilot-Quantum: L2/L3-level middleware like Covalent, Qiskit Serverless, CUDA-Q, and Pilot-Quantum tackle deployment and scaling challenges of hybrid applications. Built on tools like CUDA, MPI, Dask, and Ray, they integrate with high-level (e. g., Qiskit, Cirq) and low-level software and tools (e. g., quantum transpilers and hybrid runtimes).
L1: At the $L1$ resource layer, Qiskit Serverless, designed as a managed serverless cloud service, offers limited high-level control, e.g.., the precise number of cores, memory, and GPUs cannot be specified. It integrates well with resource-specific compilers and runtimes. Covalent also provides a managed serverless service but lacks quantum-specific integrations. Both frameworks support deployment in dedicated compute environments, typically requiring manual setup with tools like Terraform. CUDA-Q does not include a cloud or cluster resource manager and assumes ownership of resources.
Summary and Limitations: Each of the above systems targets different scenarios and quantum-HPC integration modes: Qiskit Serverless is workload-centric, Covalent focuses on workflows, CUDA-Q aims for tighter integration (HPC-for-Quantum), while the others prioritize scaling loosely coupled workloads (Quantum-x-HPC modes). Resource, workload, and task management are secondary for most systems. Pilot-Quantum addresses this gap by enabling seamless integration with heterogeneous resources, simplifying deployment, and supporting flexible resource use at the application level, crucial for diverse workloads [19].
However, Pilot-Quantum’s current implementation has limitations: it manages task dependencies at the application level, lacks high-level representations like DAGs, and restricts QPU access through classical task integration and PennyLane/Qiskit abstractions. In future work, we plan to expose QPU cloud environments through the pilot system, enabling interoperable tasks using quantum-specific languages and frameworks like OpenQASM and QIR.
V-B Conclusion and Future Work
Pilot-Quantum is middleware that enhances hybrid quantum-classical workload management on heterogeneous infrastructure. Based on analysis of execution patterns and motifs [19], it optimizes resource allocation and task execution for efficient use of heterogeneous resources. Pilot-Quantum is designed to handle complex hybrid applications with loosely coupled parallelism at our conceptual architecture’s workload layer ($L3$). In addition, it can integrate with tightly coupled application components, e. g., MPI/CUDA-based distributed state vector simulations. Our experimental results demonstrate Pilot-Quantum capabilities with significant improvements in task execution times, e. g., a 3.5×3.5$\times$3.5 × speedup for circuit cutting and a 15×15$\times$15 × speedup for QML. Pilot-Quantum will be a critical tool for quantum-HPC workload and task management, addressing resource and application heterogeneity as quantum computing evolves.
In the future, we plan to support Pilot-Quantum with higher-level frameworks like Covalent and enable direct quantum resource access via OpenQASM and QIR-based tasks. We will also investigate scheduling algorithms for heterogeneous resource task assignments. A key step will be developing Q-Dreamer, a predictive tool for estimating optimal resources for specific workloads and dynamically scheduling tasks based on real-time availability and system performance.
Acknowledgment
This research used resources from the Oak Ridge Leadership Computing Facility, a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725 (Project CSC595). Additionally, the work received support from NERSC (ERCAP0032477/ERCAP0029512: ”Characterization and Middleware for Hybrid Quantum-HPC Applications”) and the Bavarian State Ministry of Economic Affairs (BenchQC project, Grant DIK-0425/03). The authors used OpenAI’s and Anthropic’s language-generation models to generate parts of the text, which were subsequently reviewed, edited, and revised.

### v3

@@
-Pilot-Quantum: A Quantum-HPC Middleware for Resource, Workload and Task Management
+Pilot-Quantum: A Middleware for Quantum-HPC Resource, Workload and Task Management
 Abstract
-As quantum hardware continues to scale, managing the heterogeneity of resources and applications – spanning diverse quantum and classical hardware and software frameworks – becomes increasingly critical.
+As quantum hardware advances, integrating quantum processing units (QPUs) into HPC environments and managing diverse infrastructure and software stacks becomes increasingly essential.
 Pilot-Quantum addresses these challenges as a middleware designed to provide unified application-level management of resources and workloads across hybrid quantum-classical environments.
@@
 It implements the Pilot Abstraction conceptual model, originally developed for HPC, to manage resources, workloads, and tasks.
-It is designed for quantum applications that rely on task parallelism, including: (i) Hybrid algorithms, such as variational approaches, and (ii) Circuit cutting systems, used to partition and execute large quantum circuits.
-Pilot-Quantum facilitates seamless integration of quantum processing units (QPUs), classical CPUs, and GPUs, while supporting high-level programming frameworks like Qiskit and Pennylane.
-This enables users to design and execute hybrid workflows across diverse computing resources efficiently.
-The capabilities of Pilot-Quantum are demonstrated through mini-applications – simplified yet representative kernels focusing on critical performance bottlenecks.
-We present several mini-apps, including circuit execution across hardware and simulator platforms (e.
-g., IBM’s Eagle QPU), distributed state vector simulation, circuit cutting, and quantum machine learning workflows, demonstrating significant scale (e.
-g., a 41-qubit simulation on 256 GPUs) and speedups (e.
-g., 15x for QML, 3.5x for circuit cutting).
+It is designed for quantum applications that rely on task parallelism, including (i) Hybrid algorithms, such as variational approaches, and (ii) Circuit cutting systems, used to partition and execute large quantum circuits.
+Pilot-Quantum facilitates seamless integration of QPUs, classical CPUs, and GPUs, while supporting high-level programming frameworks like Qiskit and Pennylane.
+This enables users to efficiently design and execute hybrid workflows across diverse computing resources.
+The capabilities of Pilot-Quantum are demonstrated through mini-apps – simplified yet representative kernels focusing on critical performance bottlenecks.
+We demonstrate the capabilities of Pilot-Quantum through multiple mini-apps, including different circuit execution (e.g., using IBM’s Eagle QPU and simulators), circuit-cutting, and quantum machine learning scenarios.
 Index Terms:
@@
 I Introduction
-Motivation: Quantum computing promises advancements in computational challenges across science and industry, with applications in chemistry, simulations, optimization, and machine learning [1].
-Quantum algorithms leverage quantum information principles to outperform classical methods with fewer computational steps [2, 3].
-While Noisy Intermediate-Scale Quantum (NISQ) hardware [4] has progressed, further improvements in qubit fidelity [5], error correction [6], and software are needed for a practical advantage [7].
-Quantum-HPC Integration: Quantum computing complements HPC by addressing computational challenges currently intractable.
-HPC supports quantum research by enabling scalable classical simulations essential for many hybrid quantum-classical algorithms.
-Tighter integration of classical and quantum resources will become increasingly important [8, 9, 10, 11, 12].
-This integration presents challenges across several levels: (i) Hardware Integration: Efficiently combining heterogeneous quantum systems (e.g., superconducting and ion-trap qubits) with classical computing resources (e.
+Motivation: Quantum computing promises to advance specific science and industry computational challenges, from chemistry, optimization and simulations to machine learning [1], by theoretically outperforming classical methods by requiring fewer computational steps [2, 3].
+While quantum hardware has progressed, e.
+g., concerning qubit fidelity [4] and error correction [5], further improvements are needed for a practical quantum advantage [6].
+Quantum-HPC Integration: To realize the benefits of quantum computing, tighter integration of classical and quantum resources will become increasingly important [7, 8, 9, 10, 11].
+Such integration presents challenges across several levels: (i) Hardware Integration: Efficiently combining heterogeneous quantum systems (e.g., superconducting and ion-trap qubits) with classical computing resources (e.
 g., GPUs and CPUs) to create cohesive hybrid systems.
@@
 Achieving this integration is crucial for advancing quantum computing and unlocking its full potential in solving complex computational problems.
-Challenges: In particular, software tools [13, 14, 15] and middleware [16, 17, 18] will become critical for hardware development.
-The fragmented quantum software stack, still in its infancy, requires integration across multiple layers – from qubit control and circuit execution to application workflows.
-NISQ and FTQC algorithms rely heavily on quantum-classical integration [19, 20].
-With maturing algorithms, end-to-end application workflows integrating data acquisition, pre-processing, and post-processing will gain importance [21].
-Resource, workload, and task management are critical as these workflows use heterogeneous resources.
-Coordinating quantum and classical resources, managing tasks across diverse infrastructures, and integrating multiple programming paradigms complicate quantum application scaling.
-Quantum-HPC Middleware: We propose Pilot-Quantum, a middleware for resource, workload, and task management in quantum-HPC systems.
-It is based on the Pilot-Abstraction [22, 23] and a conceptual quantum-HPC middleware architecture [20], designed from an analysis of quantum-HPC execution patterns (motifs [19]).
-The pilot abstraction uses placeholder jobs to reserve and manage resources across heterogeneous infrastructures (e.
-g., HPC and clouds).
-It supports loosely coupled parallelism in quantum applications, e.
-g., variational algorithms and circuit cutting.
-Pilot-Quantum enables resource management and scheduling at the application level.
-It integrates seamlessly with quantum software ecosystems, supporting both high-level frameworks like Qiskit and Pennylane and low-level frameworks like cuQuantum for hybrid workloads [19].
-With application-level scheduling, applications can optimize resource allocation, essential for costly QPUs, enabling efficient resource use for diverse classical and quantum task-based workloads.
+Challenges: Resource, workload, and task management in classical workflow systems face significant challenges arising from heterogeneity, application and infrastructure dynamism, complex integration needs, and software fragmentation [12].
+While quantum-HPC systems share these difficulties, they substantially amplify them due to additional complexities unique to quantum computing.
+For example, leveraging quantum-HPC systems fully requires integrating multiple new types of infrastructure and support for new parallelism and integration patterns, e.g., circuit-level parallelism across multiple QPUs, tightly-coupled GPU-QPU integration for error correction, and loosely-coupled integration of variational quantum algorithms (VQA).
+Additional new task constraints, including qubit counts, fidelity thresholds, and classical-quantum resource co-location, must also be considered for scheduling and execution.
+Further, the rapidly evolving and often fragmented quantum hardware, software libraries [13, 14, 15], and middleware [16, 17, 18] landscape makes integration and scaling difficult.
+Quantum-HPC Middleware: Modern workflow systems are increasingly shifting from monolithic designs to modular architectures that separate concerns across distinct layers [12].
+Following this principle, we introduce Pilot-Quantum, a modular middleware for resource, workload, and task management that provides an application-level framework for orchestrating hybrid quantum-classical workloads across CPUs, GPUs, and QPUs.
+Pilot-Quantum addresses the key challenges of hybrid execution through multi-level scheduling, enabling adaptive resource allocation tailored to application needs and efficient use of costly and heterogeneous resources.
+Pilot-Quantum builds upon the Pilot-Abstraction [19, 20], a conceptual model for quantum-HPC middleware [21] and an analysis of execution motifs [22].
+It supports integration with high-level frameworks like Qiskit [14, 23] and PennyLane [15] and low-level runtimes like cuQuantum [24].
+This design enables scalable, efficient execution of complex hybrid workflows in rapidly evolving quantum-HPC environments.
 This paper is structured as follows: We explore quantum software and middleware state-of-the-art in section II.
-The primary contribution is Pilot-Quantum, an open-source middleware addressing resource, workload, and task management challenges is introduced in section III.
-Section IV demonstrates its capabilities by extending quantum mini apps [19] to new motifs, including distributed state vector simulation, circuit cutting, and quantum machine learning (QML).
+The primary contribution is Pilot-Quantum, an open-source middleware [25] addressing resource, workload, and task management challenges, introduced in section III.
+Section IV demonstrates its capabilities by extending quantum mini apps [22] introducing, e.g., a circuit cutting and quantum machine learning (QML) mini-apps.
 We conclude with a qualitative comparison of quantum middleware systems with Pilot-Quantum and future work in section V.
 II Background and Related Work
-This section analyzes quantum-classical software and middleware using a conceptual architecture model [20] and quantum-HPC integration modes and execution motifs [19].
+This section analyzes quantum-classical software and middleware using a conceptual architecture model [21], quantum-HPC integration modes and execution motifs [22].
 II-A Conceptual Middleware Architecture
-The conceptual quantum-HPC middleware architecture [20] consists of four distinct layers, denoted as $L1$, $L2$, $L3$, $L4$: $L4$ is the workflow layer, which encapsulates application semantics and logical dependencies between the workflow stages and offers a high-level abstraction for developing the quantum-HPC workflows.
+In [21], we define a conceptual architecture for a high-level system design of quantum-HPC middleware systems that explains the main components and their interactions.
+It consists of four distinct layers, denoted as $L1$, $L2$, $L3$, $L4$.
+$L4$ is the workflow layer, which encapsulates application semantics and logical dependencies between the workflow stages and offers a high-level abstraction for developing the quantum-HPC workflows.
 The workload layer $L3$ manages the workload, a set of interdependent, quantum, and classical tasks emitted by $L4$.
@@
 The $L2$ task layer performs the operational management of task execution on a single resource.
-Any runtime operations, i.e., task scheduling, assignment, execution, and monitoring, are managed at $L2$.
+Any runtime operations, i.
+e., task scheduling, assignment, execution, and monitoring, are managed at $L2$.
 Finally, the resource layer $L1$ describes how a system abstracts and manages heterogeneous classical (e.
@@
 II-B Quantum-HPC Integration Modes and Motifs
-Managing quantum-HPC workloads and determining quantum-classical task dependencies requires identifying integration modes and motifs (i.
-e., recurring execution patterns) between quantum-HPC components [19].
+Understanding common quantum-HPC integration modes and motifs (i.e., recurring execution patterns) between quantum-HPC components [22] is critical to designing effective middleware systems.
 Integration modes
@@
 Three modes are identified: HPC-for-Quantum, with strong coupling optimizing quantum-classical interactions (e.
-g., error mitigation [24], error correction [25, 26], dynamic circuits [27]); Quantum-in-HPC emphasizes loosely-coupled parallelism between classical and quantum resource found in variational quantum algorithms (VQA) [28] to accelerate HPC application; and Quantum-about-HPC, loosely coupling quantum elements into HPC applications for pre- and post-processing [29, 30].
+g., error mitigation [26], error correction [27, 28], dynamic circuits [29]); Quantum-in-HPC emphasizes loosely-coupled parallelism between classical and quantum resources found in VQA [30] to accelerate HPC applications; and Quantum-about-HPC, with loose coupling of quantum elements into HPC applications for pre- and post-processing [31, 32].
 Motifs
@@
 frameworks typically provide quantum circuit model-based abstraction for describing quantum programs.
-Examples are Qiskit [14, 31], PennyLane [15], Cirq [32], low-level frameworks like CUDA-Q and AWS Braket [33], and high-level programming libraries, such as, Qrisp [34], Silq [35], and Q# [36].
+Examples are Qiskit [14, 23], PennyLane [15], Cirq [33], low-level frameworks like CUDA-Q and AWS Braket [34], and high-level programming libraries, such as Qrisp [35], Silq [36], and Q# [37].
 At the $L4+$ levels, domain-specific programming libraries include Qiskit Optimization [38], Qiskit Nature [39], and OpenFermion [40].
 Hybrid Runtimes
-like Qiskit Runtime [40] and Braket Hybrid Jobs [41, 42] provide session and decorator abstractions, respectively, to co-allocate heterogeneous resources for hybrid quantum-classical applications.
+like Qiskit Runtime [41] and Braket Hybrid Jobs [42, 43] provide session and decorator abstractions, respectively, to co-allocate heterogeneous resources for hybrid applications.
 However, new execution patterns, such as the combination of quantum and machine learning capabilities, require new abstractions, including auto differentiation of quantum circuits (e.
@@
 Intermediate representations like OpenQASM [49] and QIR [50, 51] aim to bridge the gap between hardware platforms and high-level programming constructs.
-Transpiler and compiler are used to transform a circuit to a hardware-optimized form and are provided by full-stack frameworks like Qiskit and specialized frameworks like Tket [52], BQSKit [53], and Staq [54] transpilers.
+Transpilers and compilers are used to transform a circuit into a hardware-optimized form.
+They are provided by full-stack frameworks like Qiskit and specialized frameworks like Tket [52], BQSKit [53], and Staq [54] transpilers.
 Multi-level Resource Management:
@@
 g., CPUs, GPUs, QPUs), such resource management will result in suboptimal performance and utilization.
-Pilot-Quantum specifically addresses the need for an application-level resource management balancing system and application objectives.
+Pilot-Quantum specifically addresses the need for an application-level resource management by balancing system and application objectives.
 Middleware and System-level Runtimes
 at the task and workload layer ($L3$ and $L4$) are needed to orchestrate quantum workloads across integrated quantum-HPC systems.
-QML workloads, for instance, involve heterogeneous classical-quantum tasks for gradient computation, backpropagation, optimization, and circuit execution, hence requiring a systematic approach to workload and task management.
+QML workloads, for instance, involve heterogeneous classical-quantum tasks for gradient computation, backpropagation, optimization, and circuit execution, requiring a systematic approach to workload and task management.
 II-D Middleware for Resource, Workload and Task Management
 Many quantum middleware systems emerged recently including, QFaaS [58], Tierkreis [59], Qibo [60], Quantum Framework (QFw) [61], Quantum Brilliance [62], Rasqal [63] and Zapata’s Orquestra [64].
-We particularly investigate Qiskit Serverless [16], Covalent [17], and CUDA-Q [18] systems system, which have different design objectives and principles.
-II-D1 Qiskit Serverless
-[65, 66, 16, 67] serves as middleware for Qiskit, managing resources and workloads in hybrid quantum-classical infrastructures while supporting long-running applications and scaling across distributed resources (CPU, GPU, QPU).
+We particularly investigate Qiskit Serverless [16], Covalent [17], and CUDA-Q [18] systems.
+II-D1 Qiskit Serverless [65, 66, 16, 67]
+is a middleware for managing resources and workloads in hybrid quantum-classical infrastructures while supporting long-running applications and scaling across distributed resources (CPU, GPU, QPU).
 It extends cloud-based quantum hardware access to classical computing environments at the $L2$ and $L3$ layers, utilizing Apache Ray [68] to execute hybrid workloads.
@@
 QPU priority is available for computations within the same Qiskit Runtime $Session$, allowing for the co-location of quantum and classical tasks.
-II-D2 Covalent
-[17, 71] orchestrates distributed computing workflows across Cloud-HPC resources.
+II-D2 Covalent [17, 71]
+orchestrates distributed computing workflows across Cloud-HPC resources.
 Its SDK provides $L4$ level abstraction using electron and lattice decorators for defining tasks and workflows (i.
@@
 At $L4$, it provides a circuit-based programming model and a kernel API for offloading circuits to hardware or simulated QPUs.
-Application-level task management uses an MPI and CUDA-Runtime on the $L2$/$L1$ level.
+Application-level task management uses an MPI and CUDA-Runtime on the $L2$/$L1$ levels.
 III Pilot-Quantum Design and Architecture
-Pilot-Quantum enhances resource, workload, and task management across heterogeneous classical and quantum infrastructures for quantum applications and workflows.
+Pilot-Quantum manages resources, workloads, and tasks across heterogeneous classical and quantum infrastructures for quantum applications and workflows.
 It is an implementation of the Pilot Abstraction [19, 20], a conceptual abstraction for application-level scheduling [19, 75].
-It uses a placeholder job concept for application-level control of allocated resources, crucial for scaling heterogeneous workloads.
-Abstraction for Resource Management: Pilot-Quantum’s abstractions enable applications to effectively decompose workflows, simplify scaling, and manage diverse task-level parallelism.
-It can efficiently parallelize and execute operations like circuit ensembles or data-parallel QML training asynchronously and concurrently.
-Easy Deployment: Pilot-Quantum simplifies workflow deployment across distributed, heterogeneous infrastructures, including non-local quantum cloud environments, HPC MPI/CUDA environments for distributed state vector simulation, and loosely coupled workflow systems.
-It can also integrate with resource-specific optimization tools, such as QPU-specific transpilers.
+It uses placeholder jobs for application-level resource control, enabling flexible and adaptive scaling of resources and workloads.
+Abstraction for Resource Management: Pilot-Quantum enables applications to allocate resources (i.e., nodes, cores, GPUs, and QPUs) and execute task-based workloads on them.
+It provides application-level control over resources, enabling adaptive optimization of resource allocations to workloads and vice versa.
+This allows the application to dynamically acquire or release resources at runtime, such as optimizing the mix of classical and quantum resources for a specific workflow step.
+It supports typical workloads in quantum workflows, such as running circuit ensembles, decomposing larger circuits via circuit cutting, and handling data-parallel QML workloads.
+Easy Deployment: Pilot-Quantum simplifies workflow deployment across distributed, heterogeneous infrastructures, including non-local quantum clouds, HPC MPI/CUDA environments for distributed state vector simulation, and loosely coupled workflow systems.
+It can also integrate with resource-specific optimization tools like QPU transpilers.
 Application Performance: Pilot-Quantum uses hierarchical, multi-level resource management to align system- and application-level requirements, enabling application-level optimization of workload and resource allocation and use.
@@
 III-A Requirements and Design Objective
-Infrastructure Heterogeneity: The system must address both software and hardware heterogeneity, encompassing a range of QPUs such as superconducting qubits, ion traps, neutral atoms, and simulated QPUs utilizing MPI or CUDA.
-This requires compatibility with diverse frameworks and tools, including transpilers, compilers, and runtimes.
-Examples include various simulator types (e.
-g., accelerated state vector and tensor network simulations leveraging MPI or CUDA) and machine learning runtimes like PyTorch and JAX.
+Infrastructure Heterogeneity: The system must address software and hardware heterogeneity, encompassing a range of QPUs like superconducting qubits, ion traps, neutral atoms, and simulated QPUs using MPI or CUDA.
+This requires compatibility with diverse frameworks and tools, such as transpilers, compilers, and runtimes, including simulators (e.
+g., state vector and tensor network simulations with MPI or CUDA) and machine learning runtimes like PyTorch and JAX.
 Applications: Quantum-HPC applications exhibit diverse workloads [22, 76], as they evolve from monolithic applications to complex multi-stage workflows [77].
@@
 Resource management in quantum-classical computing is challenging due to diverse resource types and application demands.
-Dynamic resource allocation during workflow execution is crucial to address application and infrastructure dynamism when handling varied tasks and dependencies across heterogeneous platforms.
-This requires employing multi-level scheduling coordination with efficient resource allocation, workload partitioning, and task management strategies to further optimize application performance and system utilization.
-Abstractions: Integrating quantum computing into HPC workflows requires user-friendly, flexible abstractions that hide quantum complexity while supporting various parallelism forms (e.g., task parallelism for circuits, distributed state vector simulations using MPI and cuQuantum, and differentiable computing for QML).
+Dynamic allocation during workflow execution is crucial to address application and infrastructure dynamism when handling tasks and dependencies across heterogeneous platforms.
+Multi-level scheduling enables aligning application and system objectives, efficient resource allocation, workload partitioning, and task management strategies to optimize application performance and system utilization.
+Abstractions: Integrating quantum computing into HPC workflows requires flexible abstractions that hide quantum complexity while supporting various parallelism forms (e.g., task parallelism for circuits, distributed state vector simulations using MPI and cuQuantum, and differentiable computing for QML).
 These should facilitate efficient pipeline creation for end-to-end quantum applications with pre- and post-processing integrations.
-The abstractions must offer seamless interoperability with quantum and HPC infrastructures, further ensuring compatibility with frameworks like CUDA for GPU-accelerated tasks and MPI for distributed computing.
+The abstractions must offer seamless interoperability with quantum and HPC infrastructures, ensuring compatibility with frameworks like CUDA for GPU-accelerated tasks and MPI for distributed computing.
 Interoperability: Rapidly evolving quantum programming frameworks across all technology stack layers necessitate low-overhead integration with existing ones at both application and system levels.
@@
 While small applications can use hard-wired integration, larger hybrid systems need more adaptive resource management.
-Hence, hierarchical resource management capabilities are crucial for developing and operating quantum-enabled workflow systems efficiently.
+Hence, hierarchical resource management capabilities are crucial for efficiently developing and operating quantum-enabled workflow systems.
 Next, we present each layer of the system.
@@
 For (i), applications must define required resources in a pilot_description and submit it to Pilot-Quantum, which then acquires the resources.
-For (ii), applications must decompose their problems into a set of independently executable tasks, collectively forming the workload.
+For (ii), applications must decompose their problems into independently executable tasks, collectively forming the workload.
 Tasks can encapsulate units of quantum computations (e.
@@
 g., Dask and Ray.
-Tasks can be assigned: (i) directly to a specific pilot or (ii) to a group of pilots.
+Tasks can be assigned (i) directly to a specific pilot or (ii) to a group of pilots.
 For (ii), the pilot manager assigns tasks to pilots, optimizing throughput by load balancing across all pilots.
-Please, refer to [78] for API details.
+Please refer to [78] for API details.
 III-B2 L3/L2 Workload and Task Layers:
-After receiving the pilot_description, the Pilot-Manager allocates requested resources via the local resource management system, (steps 1-3 in Figure 2).
+After receiving the pilot_description, the Pilot-Manager allocates requested resources via the local resource management system (steps 1-3 in Figure 2).
 In HPC environments, a placeholder job, which starts the Pilot-Agent is queued via Slurm.
-Once active, the Pilot Agent manages the allocated resources and executes tasks either directly or through external execution engines.
+Once active, the Pilot Agent manages the allocated resources and executes tasks directly or through external execution engines.
 External execution engines are supported via plugins; currently, Pilot-Quantum supports Dask [79] and Ray [68].
@@
 Pilot-Quantum extends its capabilities with a plugin for Ray [68], a versatile distributed computing framework that facilitates seamless integration of GPUs (essential for quantum simulations) and custom resources like QPUs.
-For tightly coupled HPC tasks, such as distributed state vector simulations, the MPI and CUDA launch mechanisms are crucial.
+The MPI and CUDA launch mechanisms, such as distributed state vector simulations, are crucial for tightly coupled HPC tasks.
 The agent leverages Slurm’s srun command to efficiently distribute tasks across nodes, cores, and GPUs, ensuring optimal parallel execution considering machine architecture and network topology.
@@
 As described, Pilot-Quantum aims to abstract heterogeneous resources through its plugin mechanism for various runtimes.
-In addition, it can be integrated with various resource-specific tools, e.
+In addition, it can be integrated with resource-specific tools, e.
 g., QPU-specific transpilers.
 IV Evaluation and Results
-This section characterizes Pilot-Quantum’s capabilities using the Quantum Mini App framework [19, 76].
+This section characterizes Pilot-Quantum’s capabilities using the Quantum Mini-App framework [22, 76].
 Mini-apps are simplified versions of quantum-HPC applications, representing key characteristics and performance bottlenecks through recurring execution patterns (motifs).
@@
 IV-A Pilot-Quantum Characterization
+TABLE I: Average Throughput and Duration by Task Size
 […]
@@
 $8192$ $27.4$ $298.2$
-TABLE I: Average Throughput and Duration by Task Size
 […]
 IV-B Circuit Simulation Mini-App
-We extend the circuit simulation mini-app [19] with quantum hardware support (section IV-B1), an implementation of distributed state vector simulation (section IV-B2), and a circuit cutting motif (section IV-B3).
+We extend the quantum simulation mini-app [22] by adding and evaluating different cloud backends (incl.
+physical quantum hardware) (section IV-B1), an implementation of distributed state vector simulation (section IV-B2) and circuit cutting (section IV-B3).
 IV-B1 Circuit Execution:
@@
 Figure 3: Circuit Execution on IonQ, IBM Eagle and Qiskit Aer (CPU and GPU): Comparing the execution times for 1,024 random quantum circuits (2 to 28 qubits) on IonQ Quantum Cloud and Qiskit’s Aer simulators (CPU and GPU) using Pilot-Quantum (Ray) on one Perlmutter node.
-For IBM Eagle, we only executed eight random circuits for each qubit configuration and scaled their mean execution time to be comparable with the 1,024-circuit experiments.
-IonQ required 4,260 seconds for 28 qubits, while Aer (CPU) took 469.7 seconds, and Aer (GPU) completed it in just 49 seconds.
+We only executed eight random circuits for each qubit configuration for IBM Eagle and scaled their mean execution time to be comparable with the 1,024-circuit experiments.
+IonQ required 4,260 seconds for 28 qubits, while Aer (CPU) took 470 seconds, and Aer (GPU) completed it in just 49 seconds.
 The scaled execution time of the IBM Eagle QPU is 199 seconds.
-We evaluate various hardware and simulated QPU backends, specifically: (i) in-process simulators running on application-allocated resources, like Qiskit Aer [81], (ii) cloud simulators, and (iii) a QPU via the IBM cloud.
-For (i), we investigate Qiskit’s Aer CPU and GPU simulators.
-For (ii), we investigate IonQ’s cloud simulator [82].
-Further, we investigate IBM’s Eagle QPU hardware accessed via the cloud.
-We used Pilot-Quantum (Ray) on one Perlmutter node to manage the execution of $1024$ random quantum circuits with $2$-$28$ qubits across these backends.
+We evaluated various quantum backends, including (i) in-process simulators using Qiskit Aer [81], (ii) IonQ cloud simulators accessed via Qiskit IonQ provider [82], and (iii) IBM’s Eagle QPU via IBM cloud using Qiskit’s Runtime library.
+We deployed Pilot-Quantum (Ray) on a Perlmutter node to manage the execution of 1,024 random circuits across these backends, ranging from $2$ to $28$ qubits.
 […]
-For IBM Eagle, we executed only eight random circuits for each qubit configuration and scaled their mean execution time to be comparable with the $1,024$-circuit experiments.
+To demonstrate Pilot-Quantum’s capabilities in executing quantum circuits in parallel on physical quantum hardware, we assume that 256 QPUs are available with exclusive access (i.e., no queuing time).
+Due to quantum hardware costs and limited availability, we execute only eight random circuits instead of 1024.
+The average execution time for these circuits is then scaled by a factor of 4 ($1,024/256=4$) to estimate the performance for 256 QPUs processing 1,024 random circuits.
+Additionally, the time accounts for the pilot quantum compute overhead of 2 seconds for 1024 circuits on a single node with 256 tasks.
 In contrast to the simulation experiments, the execution time varies slightly but does not increase exponentially.
-IV-B2 Distributed State Vector Simulation:
-enables efficient simulation of large quantum circuits with higher qubit counts and depth by using multiple nodes and GPUs [83, 84].
-The distributed state vector motif [19] uses PennyLane’s lightning.gpu [85] v0.35.0.
-This device is a quantum circuit simulator implemented on NVIDIA’s cuQuantum [46], using a C++ backend and CUDA/MPI for distributed computations and data movements.
+IV-B2 Distributed State Vector Simulations
+enables efficient simulation of large quantum circuits with higher qubit counts and depth using multiple nodes and GPUs [83, 84].
+It is based on PennyLane’s lightning.gpu [85] (v0.41.0), which leverage a cuQuantum [24] for distributed simulation.
+We use $64$ nodes and $256$ GPUs.
 […]
-Figure 4: PennyLane lightning.gpu Distributed State Vector Simulations: Computing the expectation value of a 2-layer strongly entangling layered (SEL) circuit with and without gradient calculation using PennyLane’s lightning.gpu device on 256 A100 80 GB GPUs on Perlmutter.
-Figure 4 shows distributed state vector simulation of a 2-layer strongly entangling layered (SEL) circuit [86].
-We compare circuit execution with and without gradient computation, using adjoint differentiation [87] for gradients, which offers lower memory and computational overhead than alternatives like finite difference requiring multiple circuit evaluations.
-We investigated $31$-$37$ qubits on 256 GPUs, the upper qubit range limited by gradient computation overheads, without which can scale to $41$ qubits.
-Results indicate GPU resources must be balanced between simulation and other intensive tasks like ML for optimal performance.
-IV-B3 Circuit Cutting
-[88] enables running larger circuits on quantum hardware at the cost of increased sampling and classical processing overhead.
+Figure 4: PennyLane lightning.gpu Distributed State Vector Simulations: Computing the expectation value of a 2-layer strongly entangling layered (SEL) circuit with and without gradient calculation using PennyLane’s lightning.gpu device on 256 A100 40 GB GPUs on Perlmutter.
+Figure 4 shows distributed state vector simulation of a 2-layer strongly entangling layered (SEL) circuit [86] using up to $39$ qubits.
+We compare circuit execution (i) with and (ii) without gradient calculation.
+For (i), we use adjoint differentiation [87], which reduces memory and compute overhead compared to finite difference methods.
+For (i), we could run circuits of up to $34$ qubits, and for (ii), for up to $39$ qubits.
+The results underscore resource management’s importance, including careful memory partitioning for state vectors and gradient data.
+Figure 4 shows the distributed state vector simulation of a 2-layer strongly entangling layered (SEL) circuit [86], which is frequently utilized for classification tasks.
+We compare circuit execution with and without gradient computation.
+We employ adjoint differentiation [87] for gradient calculation, a method optimized for quantum simulations that offers lower memory and computational overhead compared to alternatives like finite difference, which necessitates multiple circuit evaluations.
+We investigate qubit counts from $30$ to $34$ using $256$ GPUs.
+The overhead of gradient calculation limits us to $34$ qubits on $256$ GPUs, while without gradient computation, we scale up to $39$ qubits.
+The results indicate that classical resources must be allocated appropriately between simulation and other computationally intensive tasks like ML to ensure optimal performance.
+IV-B3 Circuit Cutting [88]
+enables running larger circuits on quantum hardware at the cost of increased sampling and classical processing overhead.
 It decomposes larger circuits into smaller ones, referred to as subcircuits, by cutting wires and gates, allowing execution within hardware and simulator constraints.
-However, the number of required subcircuit executions increases exponentially with circuit size and number of cuts [89].
 The sampling results from the subcircuits are combined to reconstruct the original circuit outcome.
-This circuit-cutting motif uses Qiskit’s circuit cutting addon [90].
-It uses Qiskit’s EfficientSU2 ansatz, commonly used in many variational algorithms, comprising single qubit rotations and linear CNOT gates between adjacent qubits.
+However, the number of required subcircuit executions increases exponentially with circuit size and number cuts [89].
+This circuit-cutting motif uses Qiskit’s circuit-cutting addon [90].
+It utilizes Qiskit’s EfficientSU2 ansatz, commonly used in many variational algorithms, comprising single qubit rotations and linear CNOT gates between adjacent qubits.
 Qiskit’s expectation value reconstruction combines results from smaller sub-experiments.
-It provides options to run full circuit and/or circuit cutting simulations, enabling comparison of the error induced by circuit cutting.
+It provides options to run full circuit or circuit cutting simulations, enabling comparison of the circuit cutting error.
 It gathers comprehensive performance metrics throughout the process.
@@
 For full circuit simulation, we use 4 GPUs/node and distributed state vector simulation.
-In the following experiments, we use a $34$-qubit quantum circuit of the described ansatz.
+We use a $34$-qubit quantum circuit of the described ansatz in the following experiments.
 For circuit cutting, we utilize two configurations: (i) subcircuit size $17$ and (ii) subcircuit size $12$.
@@
 In scenario (i), with fewer subexperiments, optimal performance is achieved on a single node.
-Conversely, scenario (ii) shows the best performance with $2$ nodes, achieving a $\sim 1.5$ speedup, demonstrating that additional resources executed the 2,100 subexperiments faster.
-When comparing to the sequential runtimes on 1 core for (i) 58 sec and (ii) 304 sec, we were able to achieve a (i) 2.1×2.1$\times$2.1 × and (ii) 3.5×3.5$\times$3.5 × speedup.
-Scenario (iii), the distributed state vector simulation, outperforms both circuit-cutting scenarios when using more than $4$ nodes.
+Conversely, scenario (ii) shows the best performance with $2$ nodes, achieving a 1.5$\sim$ 1.5$\times$ 1.5 × speedup, demonstrating that additional resources executed the 2,100 subexperiments faster.
+When comparing to the sequential runtimes on one core for (i) 58 sec and (ii) 304 sec, we were able to achieve a (i) 2.1×2.1$\times$2.1 × and (ii) 3.5×3.5$\times$3.5 × speedup.
+The distributed state vector simulation (scenario (iii)) outperforms both circuit-cutting scenarios when using more than $4$ nodes.
 The best performance was achieved with 16 nodes, showing a $\times5$ speedup.
-In summary, different circuit-cutting strategies lead to varying optimal performance points, highlighting the importance of application-level scheduling, as enabled by Pilot-Quantum, for efficient resource and task management.
+In summary, different circuit-cutting strategies lead to varying optimal performance points, highlighting the importance of application-level scheduling for efficient resource and task management, as enabled by Pilot-Quantum.
 Hence, an adaptive execution strategy combining multiple simulation methods (e.
@@
 V-A Discussion
-This section qualitatively compares Pilot-Quantum with Qiskit Serverless, Covalent, and CUDA-Q (see section  II-D) by mapping each across the four architectural layers (see Figure 8 and Table II).
+This section qualitatively compares Pilot-Quantum with Qiskit Serverless, Covalent, and CUDA-Q by mapping each across the four architectural layers (see Figure 8 and Table II).
 L4: Qiskit Serverless and Pilot-Quantum use task-based models, while Covalent provides workflow-level abstraction at $L4$ layer.
@@
 g., error correction) and large-scale computational demands (e.
-g., distributed state vector simulations), and can be used in conjunction with Pilot-Quantum.
-While each framework supports workflow dynamism, i.
-e., enabling new task creation based on prior results, including quantum tasks, only Covalent and Pilot-Quantum allow adding resources at runtime.
-At the application-level, Qiskit Serverless is designed to integrate with Qiskit’s libraries, while Covalent and Pilot-Quantum are framework-agnostic.
+g., distributed state vector simulations).
+It can be used in conjunction with Pilot-Quantum.
+While each framework supports workflow dynamism, i.e., enabling new task creation based on prior results, including quantum tasks, only Covalent and Pilot-Quantum allow adding resources at runtime.
+At the application level, Qiskit Serverless is designed to integrate with Qiskit’s libraries, while Covalent and Pilot-Quantum are framework-agnostic.
 In contrast, CUDA-Q is optimized for tight coupling of quantum and classical tasks.
 L3: Both Covalent and Pilot-Quantum offer high-level control over resource allocation and task management.
-Qiskit Serverless allows application-level task management but not resource allocation, while CUDA-Q, uses a fixed HPC resource model with MPI and CUDA for task execution, lacking resource management capabilities and requiring manual integration with resource management systems like Slurm.
-L2: At the task execution layer, most frameworks use task-based systems like Dask and Ray for parallel execution across diverse resources, decoupling application from execution details (e.
-g., scheduling).
+In contrast, Qiskit Serverless allows application-level task management but not resource allocation.
+At the same time, CUDA-Q uses a fixed HPC resource model with MPI and CUDA for task execution, but it lacks resource management capabilities and requires manual integration with resource management systems like Slurm.
+L2: At the task execution layer, most frameworks use task-based systems like Dask and Ray for parallel execution across diverse resources, decoupling application from execution details (e.g., scheduling).
 They allow some application-level control of task and resource mapping for optimally utilizing accelerator and data parallelism.
 Resource mapping depends heavily on application details, such as the amount of data per task, making application-level scheduling crucial for optimal performance.
-CUDA-Q, with its runtime, directly manages kernel offloading to QPUs or accelerators.
+The CUDA-Q runtime directly manages task execution on QPUs or accelerators.
 While supporting some dynamism, CUDA-Q’s compiled nature makes it less flexible than other task execution systems.
@@
 g., quantum transpilers and hybrid runtimes).
-L1: At the $L1$ resource layer, Qiskit Serverless, designed as a managed serverless cloud service, offers limited high-level control, e.g.., the precise number of cores, memory, and GPUs cannot be specified.
+L1: At the $L1$ resource layer, Qiskit Serverless, designed as a managed serverless cloud service, offers limited high-level control, e.g., the precise number of cores, memory, and GPUs cannot be specified.
 It integrates well with resource-specific compilers and runtimes.
@@
 Resource, workload, and task management are secondary for most systems.
-Pilot-Quantum addresses this gap by enabling seamless integration with heterogeneous resources, simplifying deployment, and supporting flexible resource use at the application level, crucial for diverse workloads [19].
+Pilot-Quantum addresses this gap by enabling seamless integration with heterogeneous resources, simplifying deployment, and supporting flexible resource use at the application level, crucial for managing diverse workloads [22].
 However, Pilot-Quantum’s current implementation has limitations: it manages task dependencies at the application level, lacks high-level representations like DAGs, and restricts QPU access through classical task integration and PennyLane/Qiskit abstractions.
@@
 V-B Conclusion and Future Work
-Pilot-Quantum is middleware that enhances hybrid quantum-classical workload management on heterogeneous infrastructure.
-Based on analysis of execution patterns and motifs [19], it optimizes resource allocation and task execution for efficient use of heterogeneous resources.
+Pilot-Quantum is middleware that enhances resource, workload, and task management in quantum-HPC systems.
 Pilot-Quantum is designed to handle complex hybrid applications with loosely coupled parallelism at our conceptual architecture’s workload layer ($L3$).
@@
 g., MPI/CUDA-based distributed state vector simulations.
-Our experimental results demonstrate Pilot-Quantum capabilities with significant improvements in task execution times, e.
-g., a 3.5×3.5$\times$3.5 × speedup for circuit cutting and a 15×15$\times$15 × speedup for QML.
+Our experimental results demonstrate Pilot-Quantum capabilities with significant improvements in task execution times, e.g., a 3.5×3.5$\times$3.5 × speedup for circuit cutting and a 15×15$\times$15 × speedup for QML.
 Pilot-Quantum will be a critical tool for quantum-HPC workload and task management, addressing resource and application heterogeneity as quantum computing evolves.
 In the future, we plan to support Pilot-Quantum with higher-level frameworks like Covalent and enable direct quantum resource access via OpenQASM and QIR-based tasks.
-We will also investigate scheduling algorithms for heterogeneous resource task assignments.
-A key step will be developing Q-Dreamer, a predictive tool for estimating optimal resources for specific workloads and dynamically scheduling tasks based on real-time availability and system performance.
+We also aim to investigate scheduling algorithms for heterogeneous resource task assignments.
+As a future work, a key step will be developing Q-Dreamer, a predictive tool for estimating optimal resources for specific workloads and dynamically scheduling tasks based on real-time availability and system performance.
 Acknowledgment